# A Configuration-First Framework for Reproducible, Low-Code Localization

### v2

@@ IX Conclusion
We set out to reduce the practical complexity of ML-based radio localization by making reproducibility the default, lowering routine coding effort, and supporting easy extensibility. LOCALIZE delivers this with a configuration-first design: experiments are declared in human-readable files, a workflow orchestrator enforces a consistent execution, and an artifact layer captures everything needed for exact reruns and fair comparison. Ready-to-run datasets and templates further reduce non-central setup so iteration stays fast.
@@
While LOCALIZE is implemented for radio localization, its core principles offer a generalizable blueprint for building reproducible, low-code research frameworks in other scientific fields. Our configuration-first design, strict separation of concerns between experimental stages, and built-in artifact versioning are software engineering patterns that can be adapted to any domain struggling with the complexity of ML pipelines. This work can be seen as a case study in creating specialized, ’reproducible-by-design’ development environments that bridge the gap between ease of use and the rigorous methodology demanded by modern computational science.
Overall, the configuration-first approach makes localization experimentation practical, accessible, and reproducible without giving up advanced use. Future work will aim to lower any performance overheads, expand the set of preconfigured datasets, further improve and standardize extensibility, and enrich reporting and diagnostics.
Acknowledgments
This work was supported by the Slovenian Research Agency under grant P2-0016. We would like to thank Dr. Gregor Cerar for helping with the development of the early version of the software prototype.

### v3

@@ IX Conclusion
This paper addressed a core software engineering challenge in modern machine-learning research systems: how to enforce reproducibility as a default architectural property while preserving low authoring effort and extensibility. Motivated by the increasing reliance on stochastic ML components in critical software systems, we argued that reproducibility must be embedded into the structure of experimentation environments rather than treated as a post hoc obligation. We instantiated this approach through LOCALIZE, a configuration-first framework for ML-based research and studies it on a radio localization case study. Experiments are declared in human-readable configuration files, executed via an isolated workflow orchestrator, and recorded through systematic artifact lineage. Ready-to-run datasets and experiment templates further reduce non-central setup effort, enabling rapid iteration while keeping all experimental choices explicit, versioned, and auditable.